OpenStudy (anonymous):

solve the systems of equations 16x^2+4y^2=64 x^2-4=y Please show work solve the systems of equations 16x^2+4y^2=64 x^2-4=y Please show work @Mathematics

hero (hero):

muliply the second equation by -16

OpenStudy (anonymous):

I tried that but it becomes very strange

OpenStudy (anonymous):

why not substitute x^2-4 for y in the first equation? it will get a little messy, but not unbearable.

OpenStudy (anonymous):

One sec, I'll try

OpenStudy (anonymous):

I get y=-4, one sec

OpenStudy (anonymous):

x would have to be 0 in that case. So (0,-4) is one solution. There are 2 more than i can see though.

OpenStudy (anonymous):

got one, (2,-4)

OpenStudy (anonymous):

another, (-2,-4)

OpenStudy (anonymous):

and yours is the third :)

OpenStudy (anonymous):

Is that it? 3?

OpenStudy (anonymous):

(2,-4) doesnt work though. that would mean:\[x^2-4=y\iff (2)^2-4=-4\iff 4-4=-4\]

OpenStudy (anonymous):

(0, -4) is a solution though. The other ones I see are (-2,0) and (2,0). Im not sure if there are any others.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

(0,-4) doesn't work in the long equations

OpenStudy (anonymous):

nevermind